Exercise 3.1. Let X have possible values {1,2,3,4,5} and probability mass function x 1 2 3 4 5 pX(x) 1/7 1/14 3/14 2/7 2/7 (a) Calculate P(X ≤ 3). (b) Calculate P(X < 3). (c) Calculate P(X < 4.12|X > 1.638).
Answer #1


From the given data, x 1 2 3 45 11322 P()14 14 77 a) Calculate the value of P(Xs3) P(X s3) P(X 1)+P(X 2)+ P(X-3) 7 14 14 2+1+

b) Calculate the value of P(X <3) P(X < 3) = P(X = 1)+P(X-2) 7 14 2+1 14 14 0.214286 0.2143 (Round to 4 decimal place) Hence,

c) Calculate the value of P(X <4.12| X >1.638) P(X < 4.12 X > 1.638) = P(X<4.12nx1.638 P(X 1.638) P(X22) P(X 2)+P(X 3)+P(X 4)
Answer #2
Solve the possible value of X using probability mass function S = (2, 3, 4, 5).
source: Math
answered by: Aira